# SLOPE and Designing Robust Studies for Generalization

### v1

@@ Acknowledgment
The authors gratefully acknowledge Guanhua Chen, Melody Huang, Edward Kennedy, Jae-Kwang Kim, Qingyuan Zhao, and participants of the 2024 International Conference on Statistics and Data Science, 2024 Joint Statistical Meetings, 2025 American Causal Inference Conference, and the statistics seminar at Seoul National University for their valuable feedback.
@@
The Appendix is organized as follows. Section A discusses the existence of SLOPE, including the definition of SLOPE through Hadamard differentiability and some regularity conditions deferred from the main text. Section B exemplifies the use of SLOPE in choosing a robust location estimand. Section C discusses SLOPE for other estimands, including risks, quantiles, trimmed mean, OLS coefficients, some scale parameters, Pearson correlation coefficient, and general formulas of SLOPEs for L-functionals and Z-functionals. Section D discusses transporting functionals of potential outcomes. Next, Section E and Section F supplement Section 4.2 (estimation) and Section 5 (data application) in the main text, respectively, by providing details, auxiliary results, and some deferred discussions. Section G includes a simulation study that verifies the asymptotic properties of proposed estimators. Afterwards, Sections H and I provide proof for the derivation of the SLOPE (at population level) and the estimation of the SLOPE, respectively. Finally, Section J details some extended remarks, including SLOPE with a vector valued $\psi(\cdot)$, SLOPE for other types of conditional exchangeability assumptions, the challenge of extending SLOPE to bound-based sensitivity models, and the mathematical connection between SLOPE and the marginal interventional effect.
@@ A.1 Notation
We define SLOPE through the derivative of a composite of two functionals. First, define the map from $\gamma$ to $Q_{O,X}^{\gamma}$ as $\phi:\left(\mathbb{R},l^{\infty}(\mathcal{S}_{O,X})\right)\to l^{\infty}(\mathcal{S}_{O,X})$ such that $(\gamma,Q_{O,X}^{0})\mapsto Q_{O,X}^{\gamma}$, with domain $\mathbb{D}_{\phi}=\left([-\varepsilon,\varepsilon],Q_{O,X}^{0}\right)\subset\left(\mathbb{R},l^{\infty}(\mathcal{S}_{O,X})\right)$ for some $\varepsilon>0$. Next, consider the functional $\psi$ as $l^{\infty}(\mathcal{S}_{O,X})\to\mathbb{R}$ with domain $\mathbb{D}_{\psi}\subset l^{\infty}(\mathcal{S}_{O,X})$ which contains probability distributions on $\mathcal{S}_{O,X}$. Then holding $Q_{O,X}^{0}$ fixed, SLOPE as defined in (4) is the Hadamard derivative of the composite function $\psi\circ\phi$ with respect to $\gamma$ at zero. It exists under Condition 2, the standard condition that enables the chain rule of Hadamard differentiability. For Z-functionals, SLOPE exists under Condition 1 of the main text. Part (i) ensures exchanging integration and differentiation and the existence of SLOPE as an integral; parts (ii) and (iii) are analogous to regularity conditions that ensures the existence of the IF of $\psi$ (see Section 4.1).
@@ Appendix B Using SLOPE to Choose A Robust Location Parameter
$\textnormal{Corr}_{Q_{X}}[\sigma^{2}(X),f_{P_{O\mid X}}(m_{1/2})]\leq 0,\text{ where }\textnormal{Corr}\text{ represents correlation}.$
@@
where $w_{2}$ is the weight for the more heterogeneous subgroup ($X=x_{2}$) with the higher variance $\sigma_{2}^{2}$. As shown in Lemma 2, once $\mu_{1}$ and $\sigma_{2}>\sigma_{1}$ are fixed, $w_{2}$ (as a function of $\mu_{2}-\mu_{1}$) is monotonically increasing with $\mu_{2}-\mu_{1}$, the mean difference between the two subgroups. Therefore, as $w_{2}$ increases, the SLOPE for the median assigns an increasing weight for the heterogeneous subgroup, which will eventually exceed the SLOPE for the mean. Intuitively, as $\mu_{2}$ increases, the marginal target distribution under exchangeability $Q_{O}^{0}\sim 0.2N(0,0.5^{2})+0.8N(\mu_{2},0.6^{2})$ becomes more asymmetric and less uni-modal, and thus the usual understanding on median’s robustness no longer holds. Figure 5 provides a visual illustration of different underlying marginal distributions $Q_{O}^{0}$ and how they correspond to different SLOPEs for the median and the mean.
@@ C.1 SLOPE for Expectations
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{\xi})$ $=\textnormal{E}_{Q_{X}}\left\{\textnormal{Cov}_{P_{O\mid X}}\left[O,\xi(O,X)\right]\right\}.$
@@ C.5.2 SLOPE in ANCOVA
$(\alpha_{\textnormal{a}},\tau_{\textnormal{a}},\beta_{\textnormal{a}})$ $=\arg\min_{(\textnormal{W}\alpha_{\textnormal{a}},\textnormal{W}\tau_{\textnormal{a}},\textnormal{W}\beta_{\textnormal{a}})}\textnormal{E}_{Q_{O,X}^{0}}\left[\left(Y-\textnormal{W}\alpha_{\textnormal{a}}-\textnormal{W}\tau_{\textnormal{a}}A-\textnormal{W}\beta_{\textnormal{a}}^{\intercal}L\right)^{2}\right],$ (20)
$(\alpha_{\textnormal{u}},\tau_{\textnormal{u}})$ $=\arg\min_{(\textnormal{W}\alpha_{\textnormal{u}},\textnormal{W}\tau_{\textnormal{u}})}\textnormal{E}_{Q_{O,X}^{0}}\left[\left(Y-\textnormal{W}\alpha_{\textnormal{u}}-\textnormal{W}\tau_{\textnormal{u}}A\right)^{2}\right].$ (21)
@@
$\textnormal{SLOPE}(Q_{O,X}^{0},\tau_{\textnormal{a}})$ $=\textnormal{SLOPE}(Q_{O,X}^{0},\tau_{\textnormal{u}})+\delta^{\intercal}V^{-1}\delta\textnormal{Cov}_{Q_{L,A}}[A,\sigma^{2}(L,A)]-\delta^{\intercal}V^{-1}\textnormal{Cov}_{Q_{L,A}}[Z,\sigma^{2}(L,A)],$ (23)
respectively, where $\delta=\textnormal{Cov}_{Q_{X}}[L,A]/\textnormal{Var}_{Q_{A}}[A]$ and $V=\textnormal{Cov}_{Q_{Z}}(L)-\delta\delta^{\intercal}\textnormal{Var}_{Q_{A}}(A)$.
From Theorem 8, if $\textnormal{Cov}_{Q_{L,A}}(Z,A)=0$ almost surely, e.g., in cases when the intervention $A$ has been randomized, then by definition $\delta=0$ and
@@ C.7 SLOPE for Pearson Correlation Coefficient
C.8 SLOPE for L-Functionals
C.8.1 L-Functionals and Their SLOPEs
@@
Theorem 10 (SLOPE for L-Functionals).
@@ C.8.2 Lemma 2
C.9 SLOPE for Z-Functionals
For a Z-functional defined through (2) in the main text, the corresponding SLOPE is presented in the Corollary 3 below, which is an immediate result of Theorem 3.
Corollary 3 (SLOPE for Z-Functionals).
Under Condition 3, the SLOPE for a Z-functional is
@@ D.2 Transporting Functionals of a Potential Outcome
Assumption 3 (Strong Ignorability on $P$; Rosenbaum and Rubin, (1983)).
@@ Appendix E Supplementary Materials for Estimation
we provide supplementary results and some deferred discussions for the estimation of SLOPE presented in Section 4.2 of the main text. We start with defining some notations in Section E.1. Then we discuss regularity conditions and asymptotic properties of the weighting estimator and the regression estimator in Sections E.2 and E.3, respectively. Next, we detail estimators of SLOPEs for the mean, OLS coefficients, and the median in Sections E.4 to E.6, including the weighting and regression estimators presented in the main text and estimators based on the efficient influence function of the SLOPE (if exists). Then we discuss ways of estimating the nuisance function $\omega(X)$ in Section E.7. Finally, we present a general statement of the efficient influence function for the SLOPE of scalar valued Z-functionals in Section E.8.
@@ E.2 Asymptotic Properties for the Weighting Estimator
$\eta_{2}=\left\{\textnormal{E}_{Q_{X}}\left(\textnormal{E}_{P_{O\mid X}}\left[\dot{s}(O,X,\psi)\right]\right)\right\},\text{ and }\eta_{1}=\textnormal{E}_{Q_{X}}\left(\textnormal{E}_{P_{O\mid X}}\left[\{O-\mu(X)\}s\left(O,X,\psi\right)\right]\right).$
@@
With $G^{W}$ defined as above, the asymptotic properties of the weighting estimator for SLOPE in (12) can be established under standard regularity conditions in M-estimation theory (Newey and McFadden,, 1994; Van der Vaart,, 2000). Below, Condition 7 parallels the condition in Theorem 2.6 of Newey and McFadden, (1994) for consistency, and Condition 8 parallels assumptions in Theorem 3.4 of Newey and McFadden, (1994) and Theorem 5.31 of Van der Vaart, (2000) for asymptotic normality.
Condition 7 (Regularity Conditions for Consistency for Weighting Estimator).
@@
Condition 8 (Regularity Conditions for Asymptotic Normality for Weighting Estimator).
@@
With Conditions 7 and 8, we establish the asymptotic properties of the weighting estimator.
@@
Let $\widehat{\textnormal{SLOPE}}^{\textnormal{W}}$ be the weighting estimator in (12) where $\widehat{\eta}_{j}$’s are estimated with (33) and we drop the notation in parentheses of the SLOPE for ease of communication. Suppose Condition 7 holds, then $\widehat{\textnormal{SLOPE}}^{\textnormal{W}}$ converges to SLOPE in probability. Additionally suppose Condition 8 holds, then $\sqrt{n}(\widehat{\textnormal{SLOPE}}^{\textnormal{W}}-\textnormal{SLOPE})$ converges in distribution to a normal distribution with mean zero and variance $(\eta_{2})^{2}V_{11}^{\textnormal{W}}/(\eta_{1})^{4}+V_{22}^{\textnormal{W}}/(\eta_{1})^{2}-2\eta_{2}V^{W}_{12}/(\eta_{1})^{3}$, where $V^{\textnormal{W}}$ is inverse of the derivative matrix of $\textnormal{E}\{G^{\textnormal{W}}(T_{i},O_{i},X_{i},\eta^{\textnormal{W}})\}$ with respect to $\eta^{\textnormal{W}}$, with $V^{\textnormal{W}}_{ij}$ denoting its entry at the $i$-th row and $j$-th column.
@@ E.3 Asymptotic Properties for the Regression Estimator
The regularity conditions for the regression estimator are standard and are similar to the conditions for the weighting estimator. Condition 9 parallels the condition in Theorem 2.6 of Newey and McFadden, (1994) for consistency and Condition 10 parallels assumptions in Theorem 3.4 of Newey and McFadden, (1994) and Theorem 5.31 of Van der Vaart, (2000) for asymptotic normality.
Condition 9 (Regularity Conditions for Consistency for Regression Estimator).
@@
Condition 10 (Regularity Conditions for Asymptotic Normality for Regression Estimator).
@@
Let $\widehat{\textnormal{SLOPE}}^{\textnormal{R}}$ be the regression estimator in (13) where $\eta_{j}$’s are estimated with (34) and we drop the notation in parentheses of the SLOPE for ease of communication. Suppose Condition 9 in the Supplement holds, then $\widehat{\textnormal{SLOPE}}^{\textnormal{R}}$ converges to SLOPE in probability. Additionally suppose Condition 10 in the Supplement holds, then $\sqrt{n}(\widehat{\textnormal{SLOPE}}^{\textnormal{R}}-\textnormal{SLOPE})$ converges in distribution to a Gaussian distribution with mean zero and variance $(\eta_{2})^{2}V_{11}^{\textnormal{R}}/(\eta_{1})^{4}+V_{22}^{\textnormal{R}}/(\eta_{1})^{2}-2\eta_{2}V^{\textnormal{R}}_{12}/(\eta_{1})^{3}$, where $V^{\textnormal{R}}$ is the derivative matrix of $\textnormal{E}\{G^{\textnormal{R}}(T_{i},O_{i},X_{i},\eta^{\textnormal{R}})\}$ with respect to $\eta^{\textnormal{R}}$, with $V^{\textnormal{R}}_{ij}$ denoting its entry at the $i$-th row and $j$-th column.

### v2

@@ Acknowledgment
The authors gratefully acknowledge Guanhua Chen, Melody Huang, Edward Kennedy, Jae-Kwang Kim, Qingyuan Zhao, and participants of the 2024 International Conference on Statistics and Data Science, 2024 Joint Statistical Meetings, 2025 American Causal Inference Conference, and the statistics seminar at Seoul National University for their valuable feedback. The work of Xinran Miao is supported in part by funding from American Family Funding Initiative and the Morgridge Summer Research Fellowship.
@@
The Appendix is organized as follows. Section A discusses the existence of SLOPE, including the definition of SLOPE through Hadamard differentiability and some regularity conditions deferred from the main text. Section B exemplifies the use of SLOPE in choosing a robust location estimand. Section C discusses SLOPE for other estimands, including risks, quantiles, trimmed mean, OLS coefficients, some scale parameters, Pearson correlation coefficient, and general formulas of SLOPEs for L-estimands and Z-estimands. Section D discusses transporting functionals of potential outcomes. Next, Section E and Section F supplement Section 4.2 (estimation) and Section 5 (data application) in the main text, respectively, by providing details, auxiliary results, and some deferred discussions. Section G includes a simulation study that verifies the asymptotic properties of proposed estimators. Afterwards, Sections H and I provide proof for the derivation of the SLOPE (at population level) and the estimation of the SLOPE, respectively. Finally, Section J details some extended remarks, including SLOPE with a vector valued $\psi(\cdot)$, SLOPE for other types of conditional exchangeability assumptions, the challenge of extending SLOPE to bound-based sensitivity models, and the mathematical connection between SLOPE and the marginal interventional effect.
@@ A.1 Notation
We define SLOPE through the derivative of a composite of two functionals. First, define the map from $\gamma$ to $Q_{O,X}^{\gamma}$ as $\phi:\left(\mathbb{R},l^{\infty}(\mathcal{S}_{O,X})\right)\to l^{\infty}(\mathcal{S}_{O,X})$ such that $(\gamma,Q_{O,X}^{0})\mapsto Q_{O,X}^{\gamma}$, with domain $\mathbb{D}_{\phi}=\left([-\varepsilon,\varepsilon],Q_{O,X}^{0}\right)\subset\left(\mathbb{R},l^{\infty}(\mathcal{S}_{O,X})\right)$ for some $\varepsilon>0$. Next, consider the functional $\psi$ as $l^{\infty}(\mathcal{S}_{O,X})\to\mathbb{R}$ with domain $\mathbb{D}_{\psi}\subset l^{\infty}(\mathcal{S}_{O,X})$ which contains probability distributions on $\mathcal{S}_{O,X}$. Then holding $Q_{O,X}^{0}$ fixed, SLOPE as defined in (4) is the Hadamard derivative of the composite function $\psi\circ\phi$ with respect to $\gamma$ at zero. It exists under Condition 2, the standard condition that enables the chain rule of Hadamard differentiability. For Z-estimands, SLOPE exists under Condition 1 of the main text. Part (i) ensures exchanging integration and differentiation and the existence of SLOPE as an integral; parts (ii) and (iii) are analogous to regularity conditions that ensures the existence of the IF of $\psi$ (see Section 4.1).
@@ A.3 Regularity Conditions
Condition 7 (Existence of SLOPE for Sensitivity Model Defined Through $\rho(O,X,\gamma)$).
(i) $\textnormal{E}_{Q_{O\mid X}^{\gamma}}[s(O,X,\psi(Q_{O,X}^{0}))]$ is bounded for $\gamma$ in a neighborhood of zero and
$\textnormal{E}_{Q_{O,X}^{0}}\left(s(O,X,\psi(Q_{O,X}^{0}))\left[\dot{\rho}(O,X,0)-\textnormal{E}_{P_{O\mid X}}\{\dot{\rho}(O,X,0)\mid X\}\right]\right)$
exists; (ii) $s(O,X,\cdot)$ is differentiable almost everywhere with the derivative $\dot{s}(O,X,\cdot)$; (iii) $\textnormal{E}_{Q_{O,X}^{0}}\{\dot{s}(O,X,\psi(Q_{O,X}^{0}))\}$ exists and is non-singular; (iv) $\rho(O,X,\gamma)$ is twice differentiable with respect to $\gamma$ at a neighborhood of zero with derivative $\ddot{\rho}(O,X,\gamma)$.
@@ Appendix B Using SLOPE to Choose A Robust Location Parameter
$\textnormal{Corr}_{Q_{X}}[\sigma^{2}(X),f_{P_{O\mid X}}(m_{1/2}\mid X)]\leq 0,\text{ where }\textnormal{Corr}\text{ represents correlation}.$
@@
where $w_{2}$ is the weight for the more hetereogeneous subgroup ($X=x_{2}$) with the higher variance $\sigma_{2}^{2}$. As shown in Lemma 2, once $\mu_{1}$ and $\sigma_{2}>\sigma_{1}$ are fixed, $w_{2}$ (as a function of $\mu_{2}-\mu_{1}$) is monotonically increasing with $\mu_{2}-\mu_{1}$, the mean difference between the two subgroups. Therefore, as $w_{2}$ increases, the SLOPE for the median assigns an increasing weight for the heterogeneous subgroup, which will eventually exceed the SLOPE for the mean. Intuitively, as $\mu_{2}$ increases, the marginal target distribution under exchangeability $Q_{O}^{0}\sim 0.2N(0,0.5^{2})+0.8N(\mu_{2},0.6^{2})$ becomes more asymmetric and less uni-modal, and thus the usual understanding on median’s robustness no longer holds. Figure 5 provides a visual illustration of different underlying marginal distributions $Q_{O}^{0}$ and how they correspond to different SLOPEs for the median and the mean.
@@ C.1 SLOPE for Expectations
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{\xi})$ $=\textnormal{E}_{Q_{X}}\left\{\textnormal{Cov}_{P_{O\mid X}}\left[O,\xi(O,X)\mid X\right]\right\}.$
@@ C.5.2 SLOPE in ANCOVA
$(\alpha_{\textnormal{a}},\tau_{\textnormal{a}},\beta_{\textnormal{a}})$ $=\arg\min_{(\widetilde{\alpha}_{\textnormal{a}},\widetilde{\tau}_{\textnormal{a}},\widetilde{\beta}_{\textnormal{a}})}\textnormal{E}_{Q_{O,X}^{0}}\left[\left(Y-\widetilde{\alpha}_{\textnormal{a}}-\widetilde{\tau}_{\textnormal{a}}A-\widetilde{\beta}_{\textnormal{a}}^{\intercal}L\right)^{2}\right],$ (21)
$(\alpha_{\textnormal{u}},\tau_{\textnormal{u}})$ $=\arg\min_{(\widetilde{\alpha}_{\textnormal{u}},\widetilde{\tau}_{\textnormal{u}})}\textnormal{E}_{Q_{O,X}^{0}}\left[\left(Y-\widetilde{\alpha}_{\textnormal{u}}-\widetilde{\tau}_{\textnormal{u}}A\right)^{2}\right].$ (22)
@@
$\textnormal{SLOPE}(Q_{O,X}^{0},\tau_{\textnormal{a}})$ $=\textnormal{SLOPE}(Q_{O,X}^{0},\tau_{\textnormal{u}})+\delta^{\intercal}V^{-1}\delta\textnormal{Cov}_{Q_{L,A}}[A,\sigma^{2}(L,A)]-\delta^{\intercal}V^{-1}\textnormal{Cov}_{Q_{L,A}}[L,\sigma^{2}(L,A)],$ (24)
respectively, where $\delta=\textnormal{Cov}_{Q_{X}}[L,A]/\textnormal{Var}_{Q_{A}}[A]$ and $V=\textnormal{Cov}_{Q_{L}}(L)-\delta\delta^{\intercal}\textnormal{Var}_{Q_{A}}(A)$.
From Theorem 9, if $\textnormal{Cov}_{Q_{L,A}}(L,A)=0$ almost surely, e.g., in cases when the intervention $A$ has been randomized, then by definition $\delta=0$ and
@@ C.7 SLOPE for Pearson Correlation Coefficient
C.8 SLOPE for L-Estimands
C.8.1 L-Estimands and Their SLOPEs
@@
Theorem 11 (SLOPE for L-Estimands).
@@ C.8.2 Lemma 2
C.9 SLOPE for Z-Estimands
For a Z-estimand defined through (2) in the main text, the corresponding SLOPE is presented in the Corollary 2 below, which is an immediate result of Theorem 3.
Corollary 2 (SLOPE for Z-Estimands).
Under Condition 3, the SLOPE for a Z-estimand is
@@ D.2 Transporting Functionals of a Potential Outcome
Assumption 3 (Strong Ignorability on $P$; Rosenbaum and Rubin, 1983b ).
@@ Appendix E Supplementary Materials for Estimation
we provide supplementary results and some deferred discussions for the estimation of SLOPE presented in Section 4.2 of the main text. We start with defining some notations in Section E.1. Then we discuss regularity conditions and asymptotic properties of the weighting estimator and the regression estimator in Sections E.2 and E.3, respectively. Next, we detail estimators of SLOPEs for the mean, OLS coefficients, and the median in Sections E.4 to E.6, including the weighting and regression estimators presented in the main text and estimators based on the efficient influence function of the SLOPE (if exists). Then we discuss ways of estimating the nuisance function $\omega(X)$ in Section E.7. Finally, we present a general statement of the efficient influence function for the SLOPE of scalar valued Z-estimands in Section E.8.
@@ E.2 Asymptotic Properties for the Weighting Estimator
$\eta_{2}=\left\{\textnormal{E}_{Q_{X}}\left(\textnormal{E}_{P_{O\mid X}}\left[\dot{s}(O,X,\psi)\mid X\right]\right)\right\},\text{ and }\eta_{1}=\textnormal{E}_{Q_{X}}\left(\textnormal{E}_{P_{O\mid X}}\left[\{O-\mu(X)\}s\left(O,X,\psi\right)\mid X\right]\right).$
@@
With $G^{W}$ defined as above, the asymptotic properties of the weighting estimator for SLOPE in (12) can be established under standard regularity conditions in M-estimation theory (Newey and McFadden,, 1994; Van der Vaart,, 2000). Below, Condition 8 parallels the condition in Theorem 2.6 of Newey and McFadden, (1994) for consistency, and Condition 9 parallels assumptions in Theorem 3.4 of Newey and McFadden, (1994) and Theorem 5.31 of Van der Vaart, (2000) for asymptotic normality.
Condition 8 (Regularity Conditions for Consistency for Weighting Estimator).
@@
Condition 9 (Regularity Conditions for Asymptotic Normality for Weighting Estimator).
@@
With Conditions 8 and 9, we establish the asymptotic properties of the weighting estimator.
@@
Let $\widehat{\textnormal{SLOPE}}^{\textnormal{W}}$ be the weighting estimator in (12) where $\widehat{\eta}_{j}$’s are estimated with (34) and we drop the notation in parentheses of the SLOPE for ease of communication. Suppose Condition 8 holds, then $\widehat{\textnormal{SLOPE}}^{\textnormal{W}}$ converges to SLOPE in probability. Additionally suppose Condition 9 holds, then $\sqrt{n}(\widehat{\textnormal{SLOPE}}^{\textnormal{W}}-\textnormal{SLOPE})$ converges in distribution to a normal distribution with mean zero and variance $(\eta_{2})^{2}V_{11}^{\textnormal{W}}/(\eta_{1})^{4}+V_{22}^{\textnormal{W}}/(\eta_{1})^{2}-2\eta_{2}V^{W}_{12}/(\eta_{1})^{3}$, where $V^{\textnormal{W}}$ is inverse of the derivative matrix of $\textnormal{E}\{G^{\textnormal{W}}(T_{i},O_{i},X_{i},\eta^{\textnormal{W}})\}$ with respect to $\eta^{\textnormal{W}}$, with $V^{\textnormal{W}}_{ij}$ denoting its entry at the $i$-th row and $j$-th column.
@@ E.3 Asymptotic Properties for the Regression Estimator
The regularity conditions for the regression estimator are standard and are similar to the conditions for the weighting estimator. Condition 10 parallels the condition in Theorem 2.6 of Newey and McFadden, (1994) for consistency and Condition 11 parallels assumptions in Theorem 3.4 of Newey and McFadden, (1994) and Theorem 5.31 of Van der Vaart, (2000) for asymptotic normality.
Condition 10 (Regularity Conditions for Consistency for Regression Estimator).
@@
Condition 11 (Regularity Conditions for Asymptotic Normality for Regression Estimator).
@@
Let $\widehat{\textnormal{SLOPE}}^{\textnormal{R}}$ be the regression estimator in (13) where $\eta_{j}$’s are estimated with (35) and we drop the notation in parentheses of the SLOPE for ease of communication. Suppose Condition 10 in the Appendix holds, then $\widehat{\textnormal{SLOPE}}^{\textnormal{R}}$ converges to SLOPE in probability. Additionally suppose Condition 11 in the Appendix holds, then $\sqrt{n}(\widehat{\textnormal{SLOPE}}^{\textnormal{R}}-\textnormal{SLOPE})$ converges in distribution to a Gaussian distribution with mean zero and variance $(\eta_{2})^{2}V_{11}^{\textnormal{R}}/(\eta_{1})^{4}+V_{22}^{\textnormal{R}}/(\eta_{1})^{2}-2\eta_{2}V^{\textnormal{R}}_{12}/(\eta_{1})^{3}$, where $V^{\textnormal{R}}$ is the derivative matrix of $\textnormal{E}\{G^{\textnormal{R}}(T_{i},O_{i},X_{i},\eta^{\textnormal{R}})\}$ with respect to $\eta^{\textnormal{R}}$, with $V^{\textnormal{R}}_{ij}$ denoting its entry at the $i$-th row and $j$-th column.